# Insect-Scale Tailless Robot with Flapping Wings: A Simple Structure and Drive for Yaw Control

### v1

@@ IV-A Simulation Setup
Boby mass $m$ $2.0*10^{-3}$ $\rm kg$
Body intertia $J_{1}$ $1.50*10^{-7}$ $\rm kg\cdot m^{2}$
@@
Wing angles $\beta$ $18020*\pi/$ 180 $\rm rad$
$\gamma$ $18060*\pi/$ 180 $\rm rad$
Length between lift center
and wing root $l$ $40.0*10^{-3}$ $\rm m$
Time constant of first-order lag $T$ $0.013$ sec
Gravitational acceleration $g$ $9.81$ $\rm kg\cdot m/sec^{2}$
TABLE II: Control Parameters
@@
Attitude $h_{x}$ $1/0.5$ $\rm 1/sec$
$h_{y}$ $1/0.5$ $\rm 1/sec$
$k_{\eta 1}$ $1/0.1$ $\rm 1/sec$
$k_{\eta 2}$ $1/0.1$ $\rm 1/sec$
$k_{\eta 3}$ $1/0.1$ $\rm 1/sec$
$\lambda_{\omega 1}$ $1/0.1$ $\rm 1/sec$
$\lambda_{\omega 2}$ $1/0.1$ $\rm 1/sec$
$\lambda_{\omega 3}$ $1/0.1$ $\rm 1/sec$
$k_{\omega 1}$ $9.50*10^{-8}$
$k_{\omega 2}$ $8.55*10^{-8}$
$k_{\omega 3}$ $1.40*10^{-7}$
$\gamma_{\omega 1}$ $7.70*10^{-6}$
$\gamma_{\omega 2}$ $6.93*10^{-6}$
$\gamma_{\omega 3}$ $1.13*10^{-5}$
$\hat{\tau}_{o,x}(t=0)$ 00 ${\rm Nm}$
$\hat{\tau}_{o,y}(t=0)$ 00 ${\rm Nm}$
$\hat{\tau}_{o,z}(t=0)$ 00 ${\rm Nm}$
@@
$k_{z}$ $6.34*10^{-1}$
$\gamma_{z}$ $2.05*10^{-4}$
$\hat{f}_{o,z}(t=0)$ 00 ${\rm N}$
TABLE III: Unknown Offsets
Symbol Value Unit
$\delta\beta$ $180{10}*\pi/$ 180 $\rm rad$
$\delta\gamma$ $180{10}*\pi/$ 180 $\rm rad$
$\delta l$ $5*10^{-3}$ $\rm m$
${\delta f}_{w}$ $mg/4/{3}*[0,-1,0,0]^{\top}$ $\rm N$
In the simulations, we consider some case studies with lift force offsets, as shown in Table III.
@@
The misalignments $\delta\beta$ and $\delta\gamma$ that occur during manufacturing were both set to 10 degrees. The error $\delta l$ in the distance $l$ from the mounting position of each wing to the center of thrust was set to 5 mm for each wing.
@@
Case 3. After adaptating for modeling error
@@
Figure 3: Simulation result without lift force offset
Figure 4: Simulation result (Case 1)
Figure 5: Simulation result (Case 2)
Figure 6: Simulation result (Case 3)
IV-B Comparison Results
First, we performed a numerical experiment without the offsets in Table III. Here, a target value of $\boldsymbol{r}=[v_{x,d}^{B},v_{y,d}^{B},v_{z,d}^{B},\psi_{d}]^{\top}=[0.5,0.5,0.5,1]^{\top}$ was set. Note that the control parameters were adjusted to obtain similar closed-loop responses from both controllers. As shown in Fig. 6, the velocity and attitude controlled by each controller converged to the targets at about $1$ s without large overshoots.
Second, Fig. 6 shows the result for Case 1 with the misalignments shown in Table III. The control performance of the LQI controller was worse than that without the offset shown in Fig. 6. The reverse response of $v_{z}$ occurred for the LQI controller. In contrast, the control performance of the proposed controller was almost the same as that without the offset shown in Fig. 6.
Third, Fig. 6 shows the result for Case 2 with the wing force voltage modeling error shown in Table III. The control performance of the LQI controller was worse than that without the offset shown in Fig. 6. The reverse responses of $v_{x}$ and $v_{z}$ occurred for the LQI controller. $v_{y}$ had an overshoot. In contrast, for the proposed controller, the response of $v_{z}$ was the same as that in Fig. 6. However, a reverse response of $v_{x}$ occurred, and $v_{y}$ had a larger overshoot than that of the LQI controller.
Finally, Fig. 6 shows the result for Case 3 using the proposed controller with the parameters adapted for modeling errors. The response of the proposed controller was neither a reverse response nor an overshoot. The result was clearly improved compared to that shown in Fig. 6.
Figure 7: Experimental system for tethered controlled flight.
Figure 8: Yaw angle
Figure 9: Altitude
Figure 10: Velocity
V EXPERIMENT
@@
V-A Experimental Setup
@@
Our experimental system is depicted in Fig. 10. This system tracks markers attached to the body using the OptiTrack Prime 17 W motion capture system and a computer (Intel Core i9-9900K, 8-core 3.6 GHz, 64 GB of RAM), which calculates the position and orientation of the robot. These values are then sent to a control computer (Intel Core i7-7700K, 4-core 4.2 GHz, 32 GB of RAM) within $3$ ms.
The proposed controller calculates the flapping amplitude, which is then multiplied by a sinusoidal wave at 115 Hz, generated by a function generator (Precision 4050B, B&K), using a multiplier (AD633, Analog Devices). The amplitude is further amplified 30 times using an amplifier (HJPZ-0.3P×3, Matsusada Precision) and then applied to the robot through enameled wires. Note that the same flapping amplitude is applied to paired wings to generate an equal force. The control circuitry, along with the battery, can be implemented in digital circuitry when mounted in the body [22].
V-B Results
@@
The results of a flight experiment with the four tilted paired-wing robot using the proposed controller are shown in Figs. 10, 10, and 10. In Fig. 10, the result of the three paired-wing robot in [21] is also shown. Fig. 11 shows the sequential shots taken during the controlled flight. Here, the target values are set to $\psi_{d}=0$, $z_{d}=0.05$, and $v_{x,d}^{B}=v_{y,d}^{B}=0$. As $v_{x,d}^{B}=v_{y,d}^{B}=0$, the target values in the absolute coordinate system are also $v_{x,d}=v_{y,d}=0$. Considering the weak yaw torque, the target angular velocity $\boldsymbol{\omega}_{d}$ of the attitude control is derived from the target roll angle $\phi_{d}$, target pitch angle $\theta_{d}$, and current yaw angle $\psi$.
@@
The yaw angle $\psi$ shown in Fig. 10 gradually deviates from the target value $\psi_{d}=0$ but converges to a final value. The amount of the yaw drift is suppressed compared to the yaw angle of the three-paired-wing robot. On the other hand, because controlling the yaw angle reduces the lift force in the vertical direction, the altitude in Fig. 10 oscillates, but it can be controlled to tend to the target value of $z_{d}=0.05$ m. Furthermore, the translational velocity in Fig.10 is mostly controlled around the target value $v_{x,d}=v_{y,d}=0$. The oscillation is caused by ignoring the infinitesimal forces $f_{body,x}$ and $f_{body,y}$.
VI CONCLUSION
@@
In this study, a flapping-wing robot with four paired tilted wings to enable yaw control was fabricated. The wings are directly driven by piezoelectric actuators without transmission, and lift control is achieved simply by changing the voltage amplitude. However, it incurred an offset in the lift force; therefore, we designed an adaptive controller to alleviate the offset problem. Numerical experiments confirm that the proposed controller shows improved control performance compared to the LQI controller by adapting to unknown lift offsets. Finally, a tethered, controlled flight was performed, and the yaw drift was suppressed by the wing tilting arrangement and the proposed controller.
@@
We reported the first takeoff of the lightest battery-powered, tailless, flapping-wing robot (2.1 g insect scale) in [22]. The robot was equipped with a low-power circuit specifically for digital duty ratio control and a sensor unit in addition to a LiPo battery; however, it was not a controlled flight. In the future, we will develop a hover-capable, battery-powered, insect-scale, tailless, flapping-wing robot weighing less than 10 g by adopting the wing tilt arrangement and the proposed controller into the system developed in [22]. Furthermore, to quickly control the yaw angle in experiments rather than in simulations, it is essential to improve yaw torque generation.
APPENDIX A: Proof of Theorem 1
@@
$V_{\omega}=\frac{1}{2}s_{\omega}^{\top}(TJ)s_{\omega}+\frac{1}{2}\tilde{\tau}_{o}^{\top}\Gamma_{\omega}^{-1}\tilde{\tau}_{o}.$ (21)
@@
APPENDIX B: Proof of Theorem 2
@@
$V_{z}=\frac{T}{2}{s_{z}}^{2}+\frac{1}{2\gamma_{z}}{\tilde{f}_{o,z}}^{2}$ (23)

### v2

@@ 4.1 Controllability analysis
Body mass $m$ $2.0*10^{-3}$ $\rm kg$
Body inertia $J_{1}$ $1.50*10^{-7}$ $\rm kg\cdot m^{2}$
@@
Wing angles $\beta$ $20*\pi/180$ $\rm rad$
$\gamma$ $60*\pi/180$ $\rm rad$
Length between lift
center and wing root $l$ $40.0*10^{-3}$ $\rm m$
Time constant of
first-order lag $T$ $0.013$ sec
Gravitational
acceleration $g$ $9.81$ $\rm kg\cdot m/sec^{2}$
Table 2: Control Parameters
@@
Attitude $h_{x},h_{y}$ $1/0.5$ $\rm 1/sec$
$k_{\eta 1},k_{\eta 2},k_{\eta 3}$ $1/0.1$ $\rm 1/sec$
$\lambda_{\omega 1},\lambda_{\omega 2},\lambda_{\omega 3}$ $1/0.1$ $\rm 1/sec$
$k_{\omega 1}$ $9.50*10^{-8}$ -
$k_{\omega 2}$ $8.55*10^{-8}$ -
$k_{\omega 3}$ $1.40*10^{-7}$ -
$\gamma_{\omega 1}$ $7.70*10^{-6}$ -
$\gamma_{\omega 2}$ $6.93*10^{-6}$ -
$\gamma_{\omega 3}$ $1.13*10^{-5}$ -
$\hat{\tau}_{o,x}$$\hat{\tau}_{o,y}$$\hat{\tau}_{o,z}(t=0)$ 0 $\rm Nm$}
@@
$k_{z}$ $6.34*10^{-1}$ -
$\gamma_{z}$ $2.05*10^{-4}$ -
$\hat{f}_{o,z}(t=0)$ 0 $\rm N$}
4.2 Simulation setup
Tables 2 and 2 present the parameters of the plant and the proposed controller, respectively, used in the numerical simulations. Note that the parameters, particularly the value of the body inertia $J$, may differ from the actual values. The weight is set at 2.0 g considering the untethered flapping-wing robot with an integrated battery (Ozaki et al., 2023). Note that the control parameters were adjusted to obtain similar closed-loop responses from both controllers.
4.3 Case studies
In the simulations, we consider some case studies with lift force offsets, as shown in Table 3.
@@ 4.3 Case studies
The misalignments $\delta\beta$ and $\delta\gamma$ that occur during manufacturing were both set to 10 degrees. The error $\delta l$ in the distance $l$ from the mounting position of each wing to lift-force center was set to 5 mm for each wing.
@@
Case 3. After adapting for modeling errors
@@
Table 3: Unknown Offsets
Symbol Value Unit
$\delta\beta$ ${10}*\pi/180$ $\rm rad$
Case 1 $\delta\gamma$ ${10}*\pi/180$ $\rm rad$
$\delta l$ $5*10^{-3}$ $\rm m$
Cases 2 and 3 ${\delta f}_{w}$ $mg/4/{3}*[0,-1,0,0]^{\top}$ $\rm N$
(a) No lift force offset
(b) Case 1
(c) Case 2
(d) Case 3
Figure 4: Step responses. The diamonds, circles, and squares represent delay, overshoot of $10\%$ or greater, and settling times, respectively.
(a) Overshoot
(b) Delay time
(c) Settling time
Figure 5: Performance metrics
4.4 Comparison results
Figs. 4 and 5 show the step responses of both controllers and the performance metrics (overshoot, delay time, and settling time) in four cases, including those without offsets. Here, a target value of $\bm{r}=[v_{x,d}^{B},v_{y,d}^{B},v_{z,d}^{B},\psi_{d}]^{\top}=[0.5,0.5,0.5,1]^{\top}$ was set, and the settling time threshold was defined as $\pm 10\%$.
Fig. 4(a) presents the simulation result without the offsets in Table 3. As shown, the velocity and attitude controlled by each controller converged to the targets in approximately $1$ s without significant overshoots. The metrics in the no-offset case, shown in Fig. 5 were almost identical.
Second, Fig. 4(b) shows the result for Case 1 with the misalignments shown in Table 3. The control performance of the LQI controller was worse than that without the offset shown in Fig. 4(a). The reverse response of $v_{z}$ occurred for the LQI controller. In contrast, the proposed controller, as shown in Fig. 6(a), estimated the unknown offset earlier, resulting in performance metrics in Fig. 5 that were nearly identical for both the no offset and Case 1.
Third, Fig. 4(c) shows the result for Case 2 with the wing force voltage modeling error shown in Table 3. The control performance of the LQI controller was worse than that without the offset shown in Fig. 4(a). The reverse responses of $v_{x}$ and $v_{z}$ occurred for the LQI controller. $v_{y}$ had an overshoot. In contrast, for the proposed controller, the response of $v_{z}$ was the same as that in Fig. 4(a). However, a reverse response of $v_{x}$ occurred, and $v_{y}$ had a larger overshoot than that of the LQI controller as shown in Fig. 5(a). Fig. 6(b) shows the offset for Case 2. The larger overshoot is due to the fact that the estimation of the offset force $f_{o,z}$ took approximately $0.7$ s.
Finally, Fig. 4(d) shows the result for Case 3 using the proposed controller with the parameters adapted for modeling errors. The response of the proposed controller was neither a reverse response nor an overshoot. Additionally, as also shown in Fig. 5, the result was clearly improved compared to that shown in Fig. 4(c). This is because, as shown in Fig. 6(b), the estimate of the offset force $f_{o,z}$ converges early in approximately $0.1$ s.
(a) Case 1
(b) Cases 2 and 3
Figure 6: Offsets and the estimates obtained from the proposed controller
5 Experiment
@@
5.1 Experimental setup
Our experimental system is depicted in Fig. 7. This system tracks markers attached to the body using the OptiTrack Prime 17W motion capture system ($1.7$ MP, $70$-degree field of view, set at $333$ fps, NaturalPoint) and a computer (Intel Core i9-9900K, $8$-core $3.6$ GHz, $64$ GB of RAM), which calculates the position and orientation of the robot. These values are then sent to a control computer (Intel Core i7-7700K, $4$-core $4.2$ GHz, $32$ GB of RAM) within $3$ ms.
The proposed controller calculates the flapping amplitude, which is then multiplied by a sinusoidal wave at 115 Hz, generated by a function generator (Precision 4050B, B&K), using a multiplier (AD633, Analog Devices). The amplitude is further amplified 30 times using an amplifier (HJPZ-0.3P×3, Matsusada Precision) and then applied to the robot through enameled wires. Note that the same flapping amplitude is applied to paired wings to generate an equal force.
Figure 7: Experimental system for tethered controlled flight
(a) Time series response
(b) Flight trajectory
Figure 8: Experimental results
@@ 5.1 Experimental setup
5.2 Results
@@
The results of a flight experiment with the four tilted paired-wing robot using the proposed controller are shown in Fig. 8. The upper graph in Fig. 8(a) shows the yaw angle of the three paired-wing robot in (Jimbo et al., 2020). Fig. 9 shows the sequential shots taken during the controlled flight. Here, the target values are set to $\psi_{d}=0$, $z_{d}=0.05$, and $v_{x,d}^{B}=v_{y,d}^{B}=0$. As $v_{x,d}^{B}=v_{y,d}^{B}=0$, the target values in the absolute coordinate system are also $v_{x,d}=v_{y,d}=0$. Considering the weak yaw torque, the target angular velocity $\bm{\omega}_{d}$ of the attitude control is derived from the target roll angle $\phi_{d}$, target pitch angle $\theta_{d}$, and current yaw angle $\psi$.
@@
The yaw angle $\psi$ shown in the upper graph of Fig. 8(a) gradually deviates from the target value $\psi_{d}=0$ but converges to a final value. The amount of the yaw drift is suppressed compared to the yaw angle of the three-paired-wing robot. On the other hand, because controlling the yaw angle reduces the lift force in the vertical direction, the altitude in the middle graph of Fig. 8(a) oscillates, but it can be controlled to tend to the target value of $z_{d}=0.05$ m. Furthermore, the translational velocity in the lower graph of Fig. 8(a) is mostly controlled around the target value $v_{x,d}=v_{y,d}=0$. The oscillation is caused by ignoring the infinitesimal forces $f_{body,x}$ and $f_{body,y}$. Fig. 8(b) shows the flight trajectory. The flight time was approximately $1.5$ s.
6 Conclusion
@@
In this study, a flapping-wing robot with four paired tilted wings was developed to enable simultaneous control of four states, including yaw angle. The controllability Gramian was derived to quantify the controllability of the four states and to analyze the control characteristics of the fabricated configuration. The wings are directly driven by piezoelectric actuators without transmission, and lift control is achieved simply by changing the voltage amplitude. However, misalignment or errors in the lift force model could cause an offset. Therefore, we designed an adaptive controller to alleviate the offset problem. Numerical experiments confirm that the proposed controller shows improved control performance compared to the LQI controller by adapting to unknown lift offsets. Finally, a tethered, controlled flight was performed, and the yaw drift was suppressed by the tilted-wing arrangement and the proposed controller.
@@
Long-duration controlled flights will require on-board power. Controlled flights of battery-powered, tailless, flapping-wing robots weighing less than $10$ g are still challenging. In (Ozaki et al., 2023), we reported the first takeoff of the lightest battery-powered, tailless, flapping-wing robot ($2.1$ g insect scale), with a lift-to-power efficiency of approximately $5$ gf/W. The robot was equipped with a low-power circuit designed for digital duty-ratio control, a sensor unit, and a LiPo battery; however, it was not a controlled flight. In the future, we will incorporate the tilted-wing arrangement and the proposed controller into the system developed in (Ozaki et al., 2023) to accomplish untethered battery-powered controlled flight. Furthermore, to extract sufficient yaw torque, it is essential to improve the generation of lift force.
Appendix A Proof of Theorem 1
@@
$V_{\omega}=s_{\omega}^{\top}(TJ)s_{\omega}/2+\tilde{\tau}_{o}^{\top}\Gamma_{\omega}^{-1}\tilde{\tau}_{o}/2.$ (22)
@@
Appendix B Proof of Theorem 2
@@
$V_{z}=T{s^{2}_{z}}/2+{\tilde{f}^{~2}_{o,z}}/(2\gamma_{z})$ (24)